# The first catalogue of candidate white dwarf-main sequence binaries in open star clusters: A new window into common envelope evolution

### v1

@@ 3.3 SVM Input Features
As discussed in Section 2.2.4 and shown in Figure 2, WD+MS binaries occupy different regimes of CMDs and colour-colour spaces than MS stars (e.g. Rebassa-Mansergas et al., 2010, 2021; Anguiano et al., 2022; Nayak et al., 2023). Thus, we train our SVM to identify WD+MS binary candidates using colours constructed from Gaia DR3, Pan-STARRS1 and 2MASS magnitudes. Specifically, our set of input features includes 10 colours: $BP-RP$, $g-r$, $g-i$, $g-z$, $r-i$, $r-z$, $i-z$, $J-K$, $g-J$, and $r-K$. While many WD+MS systems are also located in a unique region of various CMDs (appearing bluewards of the ZAMS and redwards of most isolated WDs), we do not input any absolute magnitudes when training our SVM in order to mitigate effects caused by parallax uncertainty.
@@ 3.4 Model Definition
Our SVM contains 10 input features, hence the resulting hyperplane separating the WD+MS population from the likely MS cluster stars is high-dimensional. To visualize the SVM classifier, we highlight some example 2-D parameter spaces in Figure 3. In each of these colour-colour spaces, the WD+MS binaries are clearly separated from the MS cluster stars in both the training and testing data. The SVM plane is represented as a gradient, where the colour corresponds to the probability that a system is either a WD+MS binary or a MS cluster star. It is worth remembering that the SVM classifies systems by including all input features, so a star that appears to be a WD+MS binary but is classified as a MS star likely shares properties consistent with the MS population in higher dimensions (and vice versa). As shown with a confusion matrix in Figure 4, the SVM classifies the 20% of systems reserved for testing incredibly well. In our testing data, only $\sim 1\%$ of WD+MS binaries are classified as MS stars when adopting a prediction threshold above a probability of $P>0.5$. More importantly, using the same probability threshold, zero cluster members are classified as WD+MS binaries.
@@ 3.5 Model Evaluation
Although we find that the number of misclassified stars in our testing data is low, we recognize that there are other potential factors that could impact the performance of our model when applied to the sample of cluster stars described in Section 2. Below, we perform a series of tests to probe the expected contamination and completeness of our final catalogue. Throughout, we emphasize that given only 2 post-CE systems in OCs are currently known, our primary motivation in performing this study is not to be 100% complete, but rather to identify a set of high-probability WD+MS candidates with low contamination that can subsequently be followed-up in more detail.
3.5.1 Catalogue Contamination
Figure 5: Reddened MS stars of spectral types B8V to M8V relative to the training samples of OC members from Cantat-Gaudin et al. (2020) and SDSS WD+MS binaries in Rebassa-Mansergas et al. (2010). The thin line shows the Pickles (1998) stellar atmosphere models for an extinction of $A_{V}=0$ while the thick black line depicts an extinction of $A_{V}=3$. Both lines are coloured by the effective temperatures of the stars. It is clear that reddening of MS stars moves them along or away from the known WD+MS sample.
@@ 3.5.1 Catalogue Contamination
Figure 6: Leave-one-out cross validation (LOO-CV) results for our training sample. To investigate the performance of our SVM on ‘new’ clusters, we remove one cluster from our training sample and use it as test data with our SVM (each left out cluster is labelled). Overall, the SVM classifies the cluster members as low-probability WD+MS systems, where only a few cluster members are classified as high-probability WD+MS binaries. Specifically, zero sources are classified with probabilities above $P>0.9$. Hence, we opt to use this as a metric for defining a high-probability sample of systems in this study. The OC with the highest cluster member contamination is NGC 7086 (dark purple), which is due to its broad colours extending near the known WD+MS binaries (light purple) in colour-colour spaces.
@@
Impact of Specific Cluster Properties: As described in 3.2, our training sample is composed of both cluster members from five OCs and model stars spanning a variety of spectral types. While this sample does encompass a wide range of extinctions and colours, it is possible that some OCs investigated in this study have different overall properties than the stars used in our training data. To mimic the process of classifying an unseen cluster with our SVM, we perform a leave-one-out cross validation (LOO-CV) test on our training data. Using the LeaveOneGroupOut function from scikit-learn (Pedregosa et al., 2011), we remove one cluster from our training data and run it through our SVM model, repeating this for each of the five OCs.
@@
Figure 6 highlights the results of the LOO-CV, where each labelled cluster is the one that was left out of training and used as the test data with the SVM. Overall, the computed WD+MS probabilities of these cluster members are relatively low, where some clusters (i.e. Alessi 10 and NGC 2251) have little contamination even at medium probabilities (i.e. 0% of members are classified as WD+MS binaries at $P>0.5$). Other clusters like ASCC 41, NGC 2632 and NGC 7086 yield larger misclassification rates (2.5%, 0.83% and 4.8% at $P>0.5$, respectively). However, this is not unexpected: in various parameter spaces these clusters had colours and magnitudes that were “closest” to the sample of WD+MS stars (see Figure 6). Removing them from the training sample is therefore likely to shift the location of the hyperplane which the SVM uses to separate the samples (as in Figure 3).
@@
We emphasize that we expect our final model to have lower contamination than was found based on these LOO-CV tests as: (i) we specifically chose these five clusters to span the range of observed behaviour for the full sample and (ii) in the case of NGC 7086, two-thirds of the stars misclassified with $P>$0.6 are bright MS stars with $M_{\rm{G}}<6$ mag, which would be excluded from our sample based on the cuts described in Section 2.2.3. However, even with these caveats, we note that no cluster source in these LOO-CV tests was classified with a WD+MS probability larger than $P>0.9$, indicating that contamination from MS stars is likely minimal above this probability threshold. With this as motivation, we conservatively present only high-probability candidates identified with $P>0.9$ in our catalogue (see Section 3.6).
3.5.2 Catalogue Completeness
@@
Any photometric method to identify WD+MS stars will not be complete—especially when UV photometry is not available—because for certain combinations of binary masses and ages, the MS companion will completely dominate the SED. In addition, as described above, we purposely design our SVM model to minimize contamination from non-WD+MS sources by removing a subset of the SDSS WD+MS systems (Class 2) from Rebassa-Mansergas et al. (2010) that lie close to the ZAMS in both a Gaia CMD and multi-band colour colour spaces (Figure 2). However, as a consequence, our method may be less sensitive to WD+MS binaries with more luminous (massive) MS stars.
@@
In addition, we assume that the colours of WD+MS field binaries in Rebassa-Mansergas et al. (2010) will be similar to those of WD+MS systems in clusters. However, for particularly young clusters, it is possible that low mass M-dwarfs will still be in the pre-MS phase, possibly reducing our sensitivity to those systems. However, as seen in Palla (2012), almost all $0.1\mathrm{M_{\odot}}-0.2\mathrm{M_{\odot}}$ dwarfs evolve onto the ZAMS within $\sim 50-100$Myr. Hence, while younger OCs in our sample may be slightly less complete towards systems with very low mass companions, our method should be able to identify systems with low mass dwarfs in OCs with ages $\gtrsim 100$Myr.
Figure 7: SVM probabilities for sub-samples of systems (not) used in our SVM training. As described in Section 3.2.2, we use a sample of likely MS cluster members (gray) and WD+MS binaries from Rebassa-Mansergas et al. (2010) to train our SVM. To minimize contamination, we only use ‘Class 1’ binaries (dark pink) in our model, which on average have larger separation from the ZAMS in colour-colour spaces and a Gaia CMD than the ‘Class 2’ binaries (light pink). Here, we see that while our model is less sensitive to these types of binaries, $\sim 25\%$ of the ‘Class 2’ systems are still classified with high-probability with our SVM. Hence, these types of binaries could be present in our catalogue.
@@ 3.7 Additional Vetting of Candidates
While we already remove sources with large magnitude errors ($\sigma_{m}>0.36$) from our stellar samples, there are additional systematic effects that may lead to non-physical SEDs, which can then be (incorrectly) classified as high-probability WD+MS stars by our SVM model. In particular, it is possible that in the crowded regions of some OCs, the lower angular resolution of 2MASS may lead to a situation where individual 2MASS sources contain the flux from multiple sources which are individually resolved by Gaia/Pan-STARRS1. In addition, photometry in the catalogues can be affected by saturation—either of the star itself or due to the image quality being affected by nearby bright objects. We therefore perform a more detailed inspection of the 118 high-probability candidates by examining their SEDs, the Pan-STARRS1/2MASS catalogue quality flags, and images of the region surrounding each star.
@@
Upon visual inspection of our SEDs, we notice that the photometry of many systems behaves as expected: we observe blue ($g$-band) flux upticks of varying degrees (indicative of a hot companion), along with photometry well-fitting late-type MS stars at redder wavelengths (consistent with a MS companion). However, others exhibit abnormalities. In particular, we observe a number of systems that exhibit large flux jumps between adjacent bands, especially when transitioning from optical to infrared wavelengths–potentially indicative of the cross match issue described above.
@@
Figure 8: Photometric properties of our catalogue of 52 high-probability ($P>0.9$) candidate WD+MS binaries (stars). Each system is coloured by its SVM probability. Four high-probability candidates with spectroscopy and/or a light curve discussed in Section 5 in this paper are numbered: (1) Alessi12-c1, (2) Pleiades-c1, (3) Pleiades-c2, (4) Pleiades-c3. The MS cluster training data (gray) and the Rebassa-Mansergas et al. (2010) SDSS WD+MS sample (purple) are also shown for reference. Note that the WD+MS training data used in our SVM (only Class 1 WD+MS binaries; see Section 3.2.2) is indicated in plum, whereas the full Rebassa-Mansergas et al. (2010) sample (including both Class 1 and Class 2 WD+MS systems) is indicated in light purple. While the systems in our catalogue generally overlap with the SDSS binaries, our candidates lie closer on average to the ZAMS in a Gaia CMD and some example colour-colour spaces.
@@
We use multiple methods to identify which subset of our candidates may be affected. First, using the same set of B8V to M8V Pickles models (with $30<A_{V}<$ 3) described in Section 3.5.1 as a baseline, we select stars that have non-physical $i-z$, $z-y$, $y-J$ or $J-K$ colours (where we define “non-physical” as colours beyond the distribution of the Pickles MS models). Second, we inspect the Pan-STARRS1 colour image around the system using the Aladin Lite interactive sky atlas (Bonnarel et al., 2000), over-plotting the locations of both the Gaia and 2MASS measurements relative to the star of interest. We identify multiple cases where either (i) there are multiple Gaia sources of similar brightness close to a single 2MASS source or (ii) where there is a clear red source in very close proximity to a blue star in the Pan-STARRS1 image. In total, we identify 55 systems where we are not confident that the SED fed into our SVM model is representative of a single stellar system. We remove these from our candidate sample.
@@
In addition, when investigating Pan-STARRS1 photometric quality flags, we find that nine systems exhibit large magnitude fluctuations (standard deviations $>0.3$ mag) between multiple Pan-STARRS1 measurements in either the $g$, $r$ or $z$-bands. Upon inspecting the region around each star, we find that many of these sources were located near very bright stars that are oversaturated in the Pan-STARRS1 images. Hence, we also remove these systems from our catalogue. Finally, we carefully inspect each SED for sources with Pan-STARRS1 $g$-band magnitudes brighter than $g=14$ mag, as sources above these magnitudes can be impacted by saturation. We remove two additional objects from our catalogue due to potential saturation effects.
@@ 4.1 Photometric Properties
First, we find that 14 ($\sim 27\%$) of our high-probability candidates either overlap with, or are found slightly redward of the MS cluster stars in our training sample in the Gaia CMD. We emphasize that we provided only colour information, and not absolute magnitudes, to the SVM during the model training/candidate identification. These systems may represent stars with (i) uncertain distances (thus impacting their position on the vertical axis of Figure 8), (ii) locations in regions of higher reddening or (iii) slightly more evolved MS companions. We note that the spectroscopic sample of SDSS WD+MS binaries from Rebassa-Mansergas et al. (2010) does also contain a few stars in this region.
Second, we highlight that on average, our sample of candidate WD+MS systems is closer to the ZAMS in the Gaia CMD than the SDSS WD+MS sample. In particular, $\sim 40\%$ of our high-probability candidates are located within ±$\pm 0.25$ mag (in $BP-RP$) of the ZAMS, as compared to $<20\%$ of the SDSS WD+MS sample from Rebassa-Mansergas et al. (2010). A similar trend is observed in our sample colour-colour spaces, where our candidates tend to overlap with the redder portion of the distribution of SDSS WD+MS binaries. As the location of a given WD+MS system in these parameter spaces depends on the flux ratio and temperatures of the WD and MS stars, this effect is likely caused by a combination of the true prevalence of different types of systems as well as the relative sensitivity of our method (which relies on photometry) versus that of Rebassa-Mansergas et al. (2010) (which relies on spectroscopy) for detecting different combinations of WD+MS systems.
Indeed, Rebassa-Mansergas et al. (2021) observe a similar effect when searching for WD+MS stars in the Solar Neighbourhood using Gaia. They find that their Gaia-identified systems generally contain cooler WDs when compared to the SDSS sample of Rebassa-Mansergas et al. (2010). They hypothesize this is an observational effect where Gaia is better able to detect binaries where flux contributions from cool WDs are comparable to flux contributions from cool M-dwarf companions in comparison to the spectroscopy template matching method of Rebassa-Mansergas et al. (2010). Cooler WD primaries are also found in Gaia WD+MS binaries identified by Inight et al. (2021). While spectra for each of our candidates would be required to determine effective WD temperatures, binaries with overall redder colours seem to generally be consistent with other WD+MS systems identified in Gaia.

### v2

@@ 3.3 SVM Input Features
As discussed in Section 2.2.4 and shown in Figure 2, WD+MS binaries occupy different regimes of CMDs and color-color spaces than MS stars (e.g. Rebassa-Mansergas et al., 2010, 2021; Anguiano et al., 2022; Nayak et al., 2023). Thus, we train our SVM to identify WD+MS binary candidates using colors constructed from Gaia DR3, Pan-STARRS1 and 2MASS magnitudes. Specifically, our set of input features includes 10 colors: $BP-RP$, $g-r$, $g-i$, $g-z$, $r-i$, $r-z$, $i-z$, $J-K$, $g-J$, and $r-K$. While many WD+MS systems are also located in a unique region of various CMDs (appearing bluewards of the ZAMS and redwards of most isolated WDs), we do not input any absolute magnitudes when training our SVM in order to mitigate effects caused by parallax uncertainty.
@@ 3.4 Model Definition
Our SVM contains 10 input features, hence the resulting hyperplane separating the WD+MS population from the likely MS cluster stars is high-dimensional. To visualize the SVM classifier, we highlight some example 2-D parameter spaces in Figure 3. In each of these color-color spaces, the WD+MS binaries are clearly separated from the MS cluster stars in both the training and testing data. The SVM plane is represented as a gradient, where the color corresponds to the probability that a system is either a WD+MS binary or a MS cluster star. It is worth remembering that the SVM classifies systems by including all input features, so a star that appears to be a WD+MS binary but is classified as a MS star likely shares properties consistent with the MS population in higher dimensions (and vice versa). As shown with a confusion matrix in Figure 4, the SVM classifies the 20% of systems reserved for testing incredibly well. In our testing data, only $\sim 1\%$ of WD+MS binaries are classified as MS stars when adopting a prediction threshold above a probability of $P>0.5$. More importantly, using the same probability threshold, zero cluster members are classified as WD+MS binaries.
@@ 3.5 Model Evaluation
Although we find that the number of misclassified stars in our testing data is low, we recognize that there are other potential factors that could impact the performance of our model when applied to the sample of cluster stars described in Section 2. Below, we perform a series of tests to probe the expected contamination and completeness of our final catalog. Throughout, we emphasize that given only 2 post-CE systems in OCs are currently known, our primary motivation in performing this study is not to be 100% complete, but rather to identify a set of high-probability WD+MS candidates with low contamination that can subsequently be followed-up in more detail.
3.5.1 Catalog Contamination
Figure 5: Reddened MS stars of spectral types B8V to M8V relative to the training samples of OC members from Cantat-Gaudin et al. (2020) and SDSS WD+MS binaries in Rebassa-Mansergas et al. (2010). The thin line shows the Pickles (1998) stellar atmosphere models for an extinction of $A_{V}=0$ while the thick black line depicts an extinction of $A_{V}=3$. Both lines are colored by the effective temperatures of the stars. It is clear that reddening of MS stars moves them along or away from the known WD+MS sample.
@@ 3.5.1 Catalog Contamination
Figure 6: Leave-one-out cross validation (LOO-CV) results for our training sample. To investigate the performance of our SVM on ‘new’ clusters, we remove one cluster from our training sample and use it as test data with our SVM (each left out cluster is labelled). Overall, the SVM classifies the cluster members as low-probability WD+MS systems, where only a few cluster members are classified as high-probability WD+MS binaries. Specifically, zero sources are classified with probabilities above $P>0.9$. Hence, we opt to use this as a metric for defining a high-probability sample of systems in this study. The OC with the highest cluster member contamination is NGC 7086 (dark purple), which is due to its broad colors extending near the known WD+MS binaries (light purple) in color-color spaces.
@@
Impact of Specific Cluster Properties: As described in 3.2, our training sample is composed of both cluster members from five OCs and model stars spanning a variety of spectral types. While this sample does encompass a wide range of extinctions and colors, it is possible that some OCs investigated in this study have different overall properties than the stars used in our training data. To mimic the process of classifying an unseen cluster with our SVM, we perform a leave-one-out cross validation (LOO-CV) test on our training data. Using the LeaveOneGroupOut function from scikit-learn (Pedregosa et al., 2011), we remove one cluster from our training data and run it through our SVM model, repeating this for each of the five OCs.
@@
Figure 6 highlights the results of the LOO-CV, where each labelled cluster is the one that was left out of training and used as the test data with the SVM. Overall, the computed WD+MS probabilities of these cluster members are relatively low, where some clusters (i.e. Alessi 10 and NGC 2251) have little contamination even at medium probabilities (i.e. 0% of members are classified as WD+MS binaries at $P>0.5$). Other clusters like ASCC 41, NGC 2632 and NGC 7086 yield larger misclassification rates (2.5%, 0.83% and 4.8% at $P>0.5$, respectively). However, this is not unexpected: in various parameter spaces these clusters had colors and magnitudes that were “closest” to the sample of WD+MS stars (see Figure 6). Removing them from the training sample is therefore likely to shift the location of the hyperplane which the SVM uses to separate the samples (as in Figure 3).
@@
We emphasize that we expect our final model to have lower contamination than was found based on these LOO-CV tests as: (i) we specifically chose these five clusters to span the range of observed behaviour for the full sample and (ii) in the case of NGC 7086, two-thirds of the stars misclassified with $P>$0.6 are bright MS stars with $M_{\rm{G}}<6$ mag, which would be excluded from our sample based on the cuts described in Section 2.2.3. However, even with these caveats, we note that no cluster source in these LOO-CV tests was classified with a WD+MS probability larger than $P>0.9$, indicating that contamination from MS stars is likely minimal above this probability threshold. With this as motivation, we conservatively present only high-probability candidates identified with $P>0.9$ in our catalog (see Section 3.6).
3.5.2 Catalog Completeness
@@
Any photometric method to identify WD+MS stars will not be complete—especially when UV photometry is not available—because for certain combinations of binary masses and ages, the MS companion will completely dominate the SED. In addition, as described above, we purposely design our SVM model to minimize contamination from non-WD+MS sources by removing a subset of the SDSS WD+MS systems (Class 2) from Rebassa-Mansergas et al. (2010) that lie close to the ZAMS in both a Gaia CMD and multi-band color color spaces (Figure 2). However, as a consequence, our method may be less sensitive to WD+MS binaries with more luminous (massive) MS stars.
@@
In addition, we assume that the colors of WD+MS field binaries in Rebassa-Mansergas et al. (2010) will be similar to those of WD+MS systems in clusters. However, for particularly young clusters, it is possible that low mass M-dwarfs will still be in the pre-MS phase, possibly reducing our sensitivity to those systems. However, as seen in Palla (2012), almost all $0.1\mathrm{M_{\odot}}-0.2\mathrm{M_{\odot}}$ dwarfs evolve onto the ZAMS within $\sim 50-100$Myr. Hence, while younger OCs in our sample may be slightly less complete towards systems with very low mass companions, our method should be able to identify systems with low mass dwarfs in OCs with ages $\gtrsim 100$Myr.
Figure 7: SVM probabilities for sub-samples of systems (not) used in our SVM training. As described in Section 3.2.2, we use a sample of likely MS cluster members (gray) and WD+MS binaries from Rebassa-Mansergas et al. (2010) to train our SVM. To minimize contamination, we only use ‘Class 1’ binaries (dark pink) in our model, which on average have larger separation from the ZAMS in color-color spaces and a Gaia CMD than the ‘Class 2’ binaries (light pink). Here, we see that while our model is less sensitive to these types of binaries, $\sim 25\%$ of the ‘Class 2’ systems are still classified with high-probability with our SVM. Hence, these types of binaries could be present in our catalog.
@@ 3.7 Additional Vetting of Candidates
While we already removed sources with large magnitude errors ($\sigma_{m}>0.36$) from our stellar samples, additional systematic effects could lead to non-physical SEDs, causing some sources to then be (incorrectly) classified as high-probability WD+MS stars by our SVM model. In particular, it is possible that in the crowded regions of some OCs, the lower angular resolution of 2MASS may lead to a situation where individual 2MASS sources contain the flux from multiple sources which are individually resolved by Gaia/Pan-STARRS1. In addition, photometry in the catalogs could be affected by saturation—either of the star itself or due to the image quality being affected by nearby bright objects. We therefore performed a more detailed inspection of the 118 high-probability candidates by examining their SEDs, the Pan-STARRS1/2MASS catalog quality flags, and images of the region surrounding each star.
@@
Upon visual inspection of our SEDs, we noticed that the photometry of many systems behaved as expected: we observed blue ($g$-band) flux upticks of varying degrees (indicative of a hot companion), along with photometry well-fitting late-type MS stars at redder wavelengths (consistent with a MS companion). However, others exhibited abnormalities. In particular, we observed a number of systems that exhibited large flux jumps between adjacent bands, especially when transitioning from optical to infrared wavelengths–potentially indicative of the cross match issue described above.
@@
Figure 8: Photometric properties of our catalog of 52 high-probability ($P>0.9$) candidate WD+MS binaries (stars). Each system is colored by its SVM probability. Four high-probability candidates with spectroscopy and/or a light curve discussed in Section 5 in this paper are numbered: (1) Alessi12-c1, (2) Pleiades-c1, (3) Pleiades-c2, (4) Pleiades-c3. The MS cluster training data (gray) and the Rebassa-Mansergas et al. (2010) SDSS WD+MS sample (purple) are also shown for reference. Note that the WD+MS training data used in our SVM (only Class 1 WD+MS binaries; see Section 3.2.2) is indicated in plum, whereas the full Rebassa-Mansergas et al. (2010) sample (including both Class 1 and Class 2 WD+MS systems) is indicated in light purple. While the systems in our catalog generally overlap with the SDSS binaries, our candidates lie closer on average to the ZAMS in a Gaia CMD and some example color-color spaces.
@@
We used multiple methods to identify which subset of our candidates may be affected. First, using the same set of B8V to M8V Pickles models (with $30<A_{V}<$ 3) described in Section 3.5.1 as a baseline, we selected stars that had non-physical $i-z$, $z-y$, $y-J$ or $J-K$ colors (where we define “non-physical” as colors beyond the distribution of the Pickles MS models). Second, we inspected the Pan-STARRS1 color image around the system using the Aladin Lite interactive sky atlas (Bonnarel et al., 2000), over-plotting the locations of both the Gaia and 2MASS measurements relative to the star of interest. We identified multiple cases where either (i) there were multiple Gaia sources of similar brightness close to a single 2MASS source or (ii) where there was a clear red source in very close proximity to a blue star in the Pan-STARRS1 image. In total, we identified 55 systems where we are not confident that the SED fed into our SVM model is representative of a single stellar system. We removed these from our candidate sample.
@@
In addition, when investigating Pan-STARRS1 photometric quality flags, we found that nine systems exhibited large magnitude fluctuations (standard deviations $>0.3$ mag) between multiple Pan-STARRS1 measurements in either the $g$, $r$ or $z$-bands. Upon inspecting the region around each star, we found that many of these sources were located near very bright stars that are oversaturated in the Pan-STARRS1 images. Hence, we also removed these systems from our catalog. Finally, we carefully inspected each SED for sources with Pan-STARRS1 $g$-band magnitudes brighter than $g=14$ mag, as sources above these magnitudes can be impacted by saturation. We removed two additional objects from our catalog due to potential saturation effects.
@@ 4.1 Photometric Properties
First, we find that 14 ($\sim 27\%$) of our high-probability candidates either overlap with, or are found slightly redward of the MS cluster stars in our training sample in the Gaia CMD. We emphasize that we provided only color information, and not absolute magnitudes, to the SVM during the model training/candidate identification. These systems may represent stars with (i) uncertain distances (thus impacting their position on the vertical axis of Figure 8), (ii) locations in regions of higher reddening or (iii) slightly more evolved MS companions. We note that the spectroscopic sample of SDSS WD+MS binaries from Rebassa-Mansergas et al. (2010) does also contain a few stars in this region.
Second, we highlight that on average, our sample of candidate WD+MS systems is closer to the ZAMS in the Gaia CMD than the SDSS WD+MS sample. In particular, $\sim 40\%$ of our high-probability candidates are located within ±$\pm 0.25$ mag (in $BP-RP$) of the ZAMS, as compared to $<20\%$ of the SDSS WD+MS sample from Rebassa-Mansergas et al. (2010). A similar trend is observed in our sample color-color spaces, where our candidates tend to overlap with the redder portion of the distribution of SDSS WD+MS binaries. As the location of a given WD+MS system in these parameter spaces depends on the flux ratio and temperatures of the WD and MS stars, this effect is likely caused by a combination of the true prevalence of different types of systems as well as the relative sensitivity of our method (which relies on photometry) versus that of Rebassa-Mansergas et al. (2010) (which relies on spectroscopy) for detecting different combinations of WD+MS systems.
Indeed, Rebassa-Mansergas et al. (2021) observe a similar effect when searching for WD+MS stars in the Solar Neighbourhood using Gaia. They find that their Gaia-identified systems generally contain cooler WDs when compared to the SDSS sample of Rebassa-Mansergas et al. (2010). They hypothesize this is an observational effect where Gaia is better able to detect binaries where flux contributions from cool WDs are comparable to flux contributions from cool M-dwarf companions in comparison to the spectroscopy template matching method of Rebassa-Mansergas et al. (2010). Cooler WD primaries are also found in Gaia WD+MS binaries identified by Inight et al. (2021). While spectra for each of our candidates would be required to determine effective WD temperatures, binaries with overall redder colors seem to generally be consistent with other WD+MS systems identified in Gaia.